what is the answer for
x=3y
x+2y=30

Found 2 solutions by TimothyLamb, stanbon:
Answer by TimothyLamb(4379)

x = 3y
x + 2y = 30
---
x + 2y = 30
3y + 2y = 30
5y = 30
---
y = 6
x = 18
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php

Answer by stanbon(75887)

what is the answer for
x=3y
x+2y=30
------
Substitute for "x" and solve for "y":
3y+2y = 30
5y = 30
y = 6
-----
Solve for "x":
x = 3y = 3*6 = 18
---------------------
Cheers,
Stan H.
===================